# infocolor=gray

### v2

@@ 4.3 Universal Approximation
To establish the universal approximation property of the Free-RBF-KAN architecture, we rely on several classical results from approximation theory. We first recall the Kolmogorov–Arnold representation theorem, which provides a canonical functional decomposition for multivariate continuous functions.
@@
We are now ready to state the universal approximation theorem for non-polynomial KAN (NP-KAN).
@@
Corollary 4.6 (Universal Approximation of RBF-KAN).
5 Numerical experiments
We now turn to numerical experiments to assess the practical performance of the proposed Free-RBF-KAN architecture. To ensure a fair and meaningful comparison, we begin by outlining the configurations of the baseline models. The original Kolmogorov–Arnold Network (KAN) used in our comparisons employs cubic B-spline basis functions (i.e., B-splines of order 3). All models across the KAN family, including KAN and the various RBF-KAN counterparts, are constructed with the same number of nodes and layers, providing a controlled setting for evaluating expressive power and computational efficiency. Unless otherwise noted, all RBF-KAN variants in this study, including Free-RBF-KAN, use the Gaussian kernel as their radial basis activation.
5.1 Functional approximation
Our first experiment examines the approximation of a nonsmooth function to illustrate the importance of adaptive meshing (see Fig. 1). Following Actor et al. [2025], we consider the function:
$f(x,y)=\cos(4\pi x)+\sin(\pi y)+\sin(2\pi y)+|\sin(3\pi y^{2})|$ (22)
We evaluate a compact neural network architecture with layer sizes [2, 5, 1] across serveral model variants: MLP, KAN, FreeKnots-KAN, RBF-KAN, and Free-RBF-KAN. FreeKnots-KAN [Actor et al., 2025] introduces adaptive remeshing of the B-spline grid, enabling improved approximation accuracy compared to fixed-grid KANs. All models are trained using the LBFGS optimizer for 300 epochs with a learning rate of 1 on a regression dataset of 16,384 points processed in batches of 1,024.
The results show that the RBF-KAN substaintially outperforms the standard KAN while using fewer parameters. Both FreeKnots-KAN and its RBF-based counterpart achieve similarly low test errors; however Free-RBF-KAN attains this accuracy fewer parameters, underscoring its efficiency. These findings demonstrate that replacing B-splines with RBFs with Gaussian kernels not only improves accuracy but also reduce model complexity. Overall, RBF-KAN exhibits strong capability in approximating nonsmooth functions, with the Gaussian kernel providing the best performance among RBF variants considered.
Figure 1: (Left) Analytical solution; (Middle) Prediction by Free-RBF-KAN; (Right) Error residual.
Model Layers Basis Free Grids #Param. Test MSE
MLP [2,10,10,10,1] Tanh 261 5.26e-1
KAN [2,5,1] Spline 195 3.96e-3
FreeKnots-KAN [2,5,1] Spline 307 2.57e-4
RBF-KAN [2,5,1] RBF-Gaussian 120 6.05e-4
Free-RBF-KAN [2,5,1] RBF-Gaussian 290 2.39e-4
Table 2: Performance on nonsmooth 2D function.
5.2 Spectral bias
Our next experiment investigates the spectral bias of Free-RBF-KAN. Unlike multilayer perceptrons (MLPs), which are known to exhibit spectral bias, KAN has been shown to avoid this limitation [Wang et al., 2024b]. We assess whether this property extends to RBF-KAN by conducting a spectral bias analysis using the Neural Tangent Kernel (NTK) framework, following the methodology in Wang et al. [2024b]. Detailed NTK derivation can be found in Wang et al. [2024b, Sec. 4.1] and Wang et al. [2021, Sec. 3.1]. The NTK analysis is performed on a multiscale regression task:
$f(x)=0.1\sin(50\pi x)+\sin(2\pi x),\quad x\in[0,1],$ (23)
using MLP, KAN, and RBF-KAN variants. The input domain is discretized into 100 uniform grid points. The MLP architecture consists of 4 hidden layers with 100 neurons each, while all KAN variants use 3 hidden layers with 5 neurons. KAN employs cubic B-spline activations with 20 grid points, followed by a $\tanh$ normalization layer. RBF-KAN and Free-RBF-KAN use the same grid specifications but replace the B-splines with Gaussian RBFs. All models are trained using the mean square error (MSE) loss:
$\mathcal{L}(\theta)=\frac{1}{n}\sum_{i=1}^{n}(f(x_{i};\theta)-y_{i})^{2}.$ (24)
All KAN variants successfully learn the multiscale regression problem and outperform the MLP baseline, as shown in Fig. 2 for the approximated solutions and in Fig. 3 for the training loss. Moreover, the NTK eigenvalue spectra, shown in Fig. 4, reveal that RBF-KAN maintains a broader and and less rapidly decaying NTK eigenvalue spectrum comparable to that of KAN and MLP, indicating an absence of spectral bias (Fig. 4). Free-RBF-KAN exhibits an even wider eigenvalue spectrum after 9000 training steps, suggesting that the additional flexibility provided by adaptive centroids and shape parameters promotes faster convergence.
Figure 2: The approximation of $f$ in (23) using MLP, KAN, RBFKAN, and Free-RBFKAN.
5.3 High-dimensional regression problem
In the next set of experiments, we evaluate model performance on a high-dimensional regression task using the MNIST dataset [^1] https://docs.pytorch.org/vision/main/generated/torchvision.datasets.MNIST.html. This setting is nontrivial for traditional RBF networks, as placing and optimizing RBF centroids directly in high-dimensional spaces notoriously difficult and computationally expensive. On the other hand, RBF-KAN leverages a hierarchical architecture composed of univariate RBF functions, making it naturally suited with deep structures and far more effective at mitigating the curse of dimensionality.
Figure 5: Validation loss during training on MNIST dataset.
As shown in Fig. 5, RBF-KAN exhibits progressively improved performance when adaptive grids and adaptive smoothness are incorporated. Prior work [Yu et al., 2024] has reported that KAN performs worse than MLP on the MNIST dataset. While RBF-KAN still lags behind MLP in accuracy and requires longer training time, Free-RBF-KAN achieves better accuracy than the standard KAN, and substantially narrows the performance gap to MLP. In terms of the training time, the additional flexibility of Free-RBF-KAN does not introduce noticeable overhead compared to KAN, yet it consistently yields improved accuracy. These results demonstrate that the adaptivity in both the grid points and the kernel smoothness can enhance the model performance without significantly compromising computational efficiency. Among all the KAN-based methods, RBF-KAN achieves the fastest training speed but also the lowest accuracy. As observed in prior work, all KAN variants remain less competitive than MLP in terms of parameter efficiency, accuracy, and training time. Nonetheless, Free-RBF-KAN provides clear improvements in both accuracy and efficiency compared to RBF-KAN and standard KAN.
Type Layers Basis # Param. Test loss Training time (sec.)
MLP $28\times 28,64,10]$ Tanh 509410 6.702e-02 81.58
KAN $28\times 28,64,10]$ B-Spline 762240 1.166e-01 97.95
RBF-KAN $28\times 28,64,10]$ RBF 508160 2.020e-01 82.27
Free-RBF-KAN $28\times 28,64,10]$ RBF 525120 8.789e-02 85.81
Table 3: Model performance on MNIST dataset. The Free-RBF-KAN, which is RBF-KAN with trainable centroids and smoothness parameters of RBFs, yields the best loss.
5.4 Physics-Informed Machine Learning
Our next set of experiments focuses on physics-informed machine learning. We begin by briefly reviewing the formulation of Physics-Informed Neural Networks (PINNs), which incorporate physical laws, typically expressed as partial differential equations (PDEs), into the learning process. For further details on the PINN framework and training methodology, we refer the reader to Raissi et al. [2019]. Consider a PDE defined on a domain $\Omega$, subject to governing dynamics and boundary constraints. Such a system can be written in the general form
$\mathcal{N}_{\textbf{x}}[u(\textbf{x})]$ $=f(\textbf{x}),\quad\textbf{x}\in\Omega$ (25)
$\mathcal{B}_{\textbf{x}}[u(\textbf{x})]$ $=g(\textbf{x}),\quad\textbf{x}\in\partial\Omega$
where $\mathcal{N}$ is a differential operator that governs the behavior of the unknown solution $u$, and $\mathcal{B}$ is the boundary operator. The functions $f$ and $g$ represent the source term and boundary condition, respectively. Here, $\Omega$ denotes the domain of interest, and $\partial\Omega$ its boundary. PINNs solve PDEs by approximating the solution $u$ with a neural network $\hat{u}(\textbf{x};\omega)$, where $\omega$ denotes the trainable parameters. Physical laws are incorporated into the training objective through a loss function that penalizes violations of the PDE and boundary conditions. Automatic differentiation (AD) is used to efficiently compute all required derivatives at training points. The total loss combines contributions from the PDE residual evaluated at interior collocation points and the boundary residual, and is given by
$\begin{split}\mathcal{L}(\omega)&=\frac{\lambda_{c}}{N_{c}}\sum_{i=1}^{N_{c}}\left|\mathcal{N}_{\textbf{x}}[\hat{u}(\textbf{x}_{i};\omega)]-f(\textbf{x}_{i})\right|^{2}\\ &\quad+\frac{\lambda_{b}}{N_{b}}\sum_{i=1}^{N_{b}}\left|\hat{u}(\textbf{x}_{i};\omega)-g(\textbf{x}_{i})\right|^{2}\end{split}$ (26)
where $N_{c}$ and $N_{b}$ denote the number of collocation points sampled in the interior of the domain and the boundary, respectively. The coefficients $\lambda_{c}$ and $\lambda_{b}$ weight the interior and boundary losses and, unless otherwise specified, are set to $\lambda_{c}=\lambda_{b}=1$. The sampling points are typically drawn uniformly from $\Omega$ and $\partial\Omega$. Training proceeds by minimizing $\mathcal{L}(\omega)$ via gradient-based optimization to adjust the network parameters, so that the output solution satisfies both the governing PDE and the boundary conditions.
In the remainder of this section, we demonstrate the ability of RBF-KAN on physics-informed machine learning settings. Specifically, we investigate tow representative PDE problems: a 2D heat conduction with a high frequency force term, and a 2D Helmholtz equation with smooth sinusoidal source. In both cases, the models are trained using the strong-form PINN loss defined in Eq. 26 to learn the corresponding solutions.
5.4.1 Heat Conduction in 2D
The heat conduction problem from Wang et al. [2021] has been shown to be more effectively learned by in KAN than by MLP [Wang et al., 2024b]. To evaluate whether the proposed RBF-KAN can also handle multiscale physics-informed problems, we compare RBF-KAN and Free-RBF-KAN against MLP and the original KAN. The governing heat conduction equation can be written as
$\begin{split}u_{t}&=\frac{1}{(K\pi)^{2}}u_{xx},\quad x\in[0,1],t\in[0,1]\\ u(x,0)&=\sin(K\pi x),\quad x\in[0,1]\\ u(0,t)=u(1,t)&=0,t\in[0,1]\end{split}$ (27)
and its analytical solution is given by
$u(x,t)=e^{-t}\sin(K\pi x).$ (28)
In this experiment, we set $K=50$. We randomly sample 4000 interior collocation points and 200 boundary points for each boundary segment. The MLP baseline has 4 hidden layers with width 40, while all KAN variants use an architecture with layer sizes [2,5,5,1] and 30 grid points per activation. The KAN model uses cubic B-spline activations, and the grid range of KAN variations is set as $(x_{l},x_{r})=(0,1)$, same as Wang et al. [2024b]. All models are trained using the Adam optimizer with a learning rate $10^{-3}$ and an exponential learning rate scheduler with $\gamma=0.999$ applied each epoch. Training is performed for 15,000 epochs.
@@ 5.4.1 Heat Conduction in 2D
(k) KAN
(l) Relative L2 error
Figure 6: Predicted solutions from MLP, Free-RBF-KAN, RBF-KAN, and KAN on the heat conduction problem.
The results presented in Fig. 6 show that both RBF-KAN and Free-RBF-KAN can successfully learn this high frequency problem. While RBF-KAN attains slightly lower accuracy than KAN, it trains in roughly half the time required by KAN. Incorporating adaptive grids and smoothness further improves the performance: Free-RBF-KAN achieves higher accuracy than both KAN and RBF-KAN. Although the additional flexibilities in Free-RBF-KAN increase the training time slighly, Free-RBF-KAN remains substantially faster than KAN, as shown in (Table 4).
Model Layers Basis # Param. $L^{\infty}$-loss Training time
MLP [2,40,40,40,1] Tanh 5081 1 60
Free-RBF-KAN [2,5,5,1] RBF-Gaussian 2000 2.41e-3 138
RBF-KAN [2,5,5,1] RBF-Gaussian 1280 2.78e-3 124
KAN [2,5,5,1] Spline 1400 6.52e-3 267
Table 4: Performance on 2D heat conduction problem. Grid size is set 30 for all KAN approaches. The training times are measured in seconds.
5.4.2 Helmholtz Equation in 2D
To further demonstrate the benefits of adaptive meshing, we examine physics-informed learning for a 2D Helmholtz equation with Dirichlet boundary conditions, following the setup in Wang et al. [2020]. The problem is defined as
$\begin{split}u_{xx}(x,y)+u_{yy}(x,y)+k^{2}u(x,y)&=q(x,y),\quad x,y\in\Omega\\ u(x,y)&=0\quad x,y\in\partial\Omega\end{split}$ (29)
where the forcing term is given by
$q(x,y)=$ $-(a_{1}\pi)^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y)-(a_{2}\pi)^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y)$ (30)
$+k^{2}\sin(a_{1}\pi x)\sin(a_{2}\pi y),$ (31)
and the analytical solution is
$u(x,y)=\sin(a_{1}\pi x)\sin(a_{2}\pi y),$ (32)
We set $\Omega\in[-3,3]\times[-3,3]$ and choose $a_{1}=a_{2}=k=1$. For each training epoch, we randomly sampled 4,000 collocation points in the interior and 100 boundary points on each edge of the domain. The MLP baseline has 5 hidden layers with 128 nodes each, whereas all the KAN variants have 2 hidden layers of 5 nodes per layer and 10 grids for each activation function. Training is performed using the Adam optimizer with a learning rate of $10^{-3}$.

### v3

@@ 4.3 Universal Approximation
To establish the universal approximation property of the Free-RBF-KAN architecture, we rely on several classical results from approximation theory. We begin by recalling the Kolmogorov–Arnold representation theorem, which provides a canonical decomposition of multivariate continuous functions into sums of compositions of univariate functions.
@@
We are now ready to state the universal approximation theorem for non-polynomial KAN (NP-KAN). We provide a sketch of the proof here, and a full proof in Appendix 1.
@@
Proof.
(Sketch, see Section 1 for the full proof.) By Lemma 4.1, any $f\in C([0,1]^{d})$ admits decomposition (5) with continuous univariate functions $\Phi^{(q)}$ and $\phi^{(pq)}$. The inner sum $S_{q}(x)=\sum_{p}\phi^{(pq)}(x_{p})$ has compact range on which $\Phi^{(q)}$ is uniformly continuous. Since $K$ is continuous and non-polynomial, Lemma 4.3 implies that $\mathcal{S}$ is dense in $C[a,b]$ for any compact interval. So, $\forall\varepsilon>0$, each $\phi^{(pq)}$ and $\Phi^{(q)}$ can be uniformly approximated on their respective compact domains by $\widehat{\phi}^{(pq)},\widehat{\Phi}^{(q)}\in\mathcal{S}$ with errors chosen small enough so that uniform continuity controls the composition error. Summing the resulting error bounds over $q$ yields (9). ∎
Corollary 4.6 (Universal Approximation of RBF-KAN).
5 Numerical experiments
We now turn to numerical experiments to assess the practical performance of the proposed Free-RBF-KAN architecture. To ensure a fair and meaningful comparison, we begin by outlining the configurations of the baseline models. The KAN used in our comparisons employs cubic B-spline basis functions (i.e., B-splines of order 3). All models in the KAN family, including KAN and the various RBF-KAN counterparts, are constructed with the same number of nodes and layers, providing a controlled setting for evaluating expressive power and computational efficiency. Unless otherwise noted, all RBF-KAN variants, including Free-RBF-KAN, use the Gaussian kernel as their radial basis activation.
5.1 Functional approximation
Our first experiment examines the approximation of a nonsmooth function to illustrate the importance of adaptive meshing. Following Actor et al. [2025], we consider the function defined for $(x,y)\in(0,1)^{2}$:
$f(x,y)=\cos(4\pi x)+\sin(\pi y)+\sin(2\pi y)+|\sin(3\pi y^{2})|$
We evaluate a compact neural network architecture with layer sizes [2, 5, 1] across several model variants: MLP, KAN, FreeKnots-KAN, RBF-KAN, and Free-RBF-KAN. FreeKnots-KAN [Actor et al., 2025] introduces adaptive remeshing of the B-spline grid, enabling improved approximation accuracy compared to fixed-grid KANs. All models are trained using the LBFGS optimizer for 300 epochs with a learning rate of 1 on a regression dataset of 16,384 points processed in batches of 1,024.
In Table 2, the results show that the RBF-KAN substantially outperforms the standard KAN while using fewer parameters. Both FreeKnots-KAN and its RBF-based counterpart achieve similarly low test errors; however Free-RBF-KAN attains this accuracy with fewer parameters, underscoring its efficiency. These findings demonstrate that replacing B-splines with RBFs with Gaussian kernels not only improves accuracy but also reduces model complexity. Overall, RBF-KAN exhibits strong capability in approximating nonsmooth functions, with the Gaussian kernel providing the best performance among RBF variants considered.
5.2 Spectral bias
Our next experiment investigates the spectral bias of Free-RBF-KAN. Unlike MLPs, which are known to exhibit spectral bias, KAN has been shown to avoid this limitation [Wang et al., 2024b]. We assess whether this property extends to RBF-KAN by conducting a spectral bias analysis using the Neural Tangent Kernel (NTK) framework, following the methodology in Wang et al. [2024b]. Detailed NTK derivation can be found in Wang et al. [2024b, Sec. 4.1] and Wang et al. [2021, Sec. 3.1]. The NTK analysis is performed on a multiscale regression task:
$f(x)=0.1\sin(50\pi x)+\sin(2\pi x),\quad x\in[0,1],$ (10)
using MLP, KAN, and RBF-KAN variants. The input domain is discretized into 100 uniform grid points. The MLP architecture consists of 4 hidden layers with 100 neurons each, while all KAN variants use 3 hidden layers with 5 neurons. KAN employs cubic B-spline activations with 20 grid points, followed by a tanh normalization layer. RBF-KAN and Free-RBF-KAN use the same grid specifications but replace the B-splines with Gaussian RBFs. All models are trained using the mean square error (MSE) loss. All KAN variants successfully learn the multiscale regression problem and outperform the MLP baseline, as shown in Fig. 1 for the approximated solutions and in Fig. 2 for the training loss. Moreover, the NTK eigenvalue spectra, shown in Fig. 3, reveal that RBF-KAN maintains a broader and less rapidly decaying NTK eigenvalue spectrum comparable to that of KAN and MLP, indicating an absence of spectral bias (Fig. 3). Free-RBF-KAN exhibits an even wider eigenvalue spectrum after 9000 training steps, suggesting that the additional flexibility provided by adaptive centroids and shape parameters promotes faster convergence.
Figure 1: The approximation of $f$ in (10) using MLP, KAN, RBF-KAN, and Free-RBF-KAN.
5.3 High-dimensional regression problem
Regression
Problem Model Layers Basis #Param. Test MSE Training time (sec.)
2D Nonsmooth Function MLP [2,10,10,10,1] Tanh 261 5.26e-1 162
KAN [2,5,1] B-Spline 195 3.54e-3 124
FreeKnots-KAN [2,5,1] B-Spline 307 3.54e-4 136
RBF-KAN [2,5,1] RBF-Gaussian 120 5.67e-3 122
Free-RBF-KAN [2,5,1] RBF-Gaussian 450 2.74e-4 121
Model Layers Basis # Param. Test loss Training time (sec.)
MNIST MLP $28\times 28,64,10]$ Tanh 509410 6.702e-02 81.58
KAN $28\times 28,64,10]$ B-Spline 762240 1.166e-01 97.95
RBF-KAN $28\times 28,64,10]$ RBF-Gaussian 508160 2.020e-01 82.27
Free-RBF-KAN $28\times 28,64,10]$ RBF-Gaussian 525120 8.789e-02 85.81
Physics-Informed Machine Learning
Model Layers Basis # Param. $L^{2}$-loss Training time (sec.)
2D Heat Conduction MLP [2,40,40,40,1] Tanh 5081 1 60
KAN [2,5,5,1] B-Spline 1400 6.52e-3 267
RBF-KAN [2,5,5,1] RBF-Gaussian 1280 2.78e-3 124
Free-RBF-KAN [2,5,5,1] RBF-Gaussian 2000 2.41e-3 138
2D Helmholtz MLP [2,128,128,128,1] Tanh 50049 4.15e-2 39
KAN [2,5,5,1] B-Spline 600 1.58 153
RBF-KAN [2,5,5,1] RBF-Gaussian 400 3.67e-1 49
Free-RBF-KAN [2,5,5,1] RBF-Gaussian 640 3.35e-2 62
Operator Learning
Model Layers (DeepONet Trunk) Basis # Param $L^{2}$-loss Training time (sec.)
Reaction Diffusion PDE MLP [2,[40]*4,100] Tanh 18921 2.08e-2 78
KAN [2, 4, 4, 4, 100] B-Spline 11945 6.15e-2 96
RBF-KAN [2, 4, 4, 4, 100] RBF-Gaussian 10625 3.7e-2 84
Free-RBF-KAN [2, 4, 4, 4, 100] RBF-Gaussian 11185 1.94e-2 88
Table 2: Performance benchmarks across regression, physics-informed, and operator learning tasks. We evaluate FreeKnots-KAN [Actor et al., 2025] on 2D nonsmooth function. All B-spline basis functions utilize a third-order smoothness. For operator learning tasks, the DeepONet branch architecture consists of an MLP with [100, [40]*4,100] neurons, evaluated across various trunk configurations.
In the next set of experiments, we evaluate model performance on a high-dimensional regression task using the MNIST dataset [^1] https://docs.pytorch.org/vision/main/generated/torchvision.datasets.MNIST.html. This setting is nontrivial for traditional RBF networks, as placing and optimizing RBF centroids directly in high-dimensional spaces notoriously difficult and computationally expensive. On the other hand, RBF-KAN leverages a hierarchical architecture composed of univariate RBF functions, making it naturally suited to deep structures and far more effective at mitigating the curse of dimensionality.
RBF-KAN exhibits progressively improved performance when adaptive grids and adaptive smoothness are incorporated. Prior work [Yu et al., 2024] has reported that KAN performs worse than MLP on the MNIST dataset. While RBF-KAN still lags behind MLP in accuracy and requires longer training time, Free-RBF-KAN achieves better accuracy than the standard KAN, and substantially narrows the performance gap to MLP. In terms of the training time, the additional flexibility of Free-RBF-KAN does not introduce noticeable overhead compared to KAN, yet it consistently yields improved accuracy. These results demonstrate that the adaptivity in both the grid points and the kernel smoothness can enhance the model performance without significantly compromising computational efficiency. Among all the KAN-based methods, RBF-KAN achieves the fastest training speed but also the lowest accuracy. As observed in prior work Yu et al. [2024], all KAN variants remain less competitive than MLP in terms of parameter efficiency, accuracy, and training time. Nonetheless, Free-RBF-KAN provides clear improvements in both accuracy and efficiency compared to RBF-KAN and standard KAN.
5.4 Physics-Informed Machine Learning
We evaluate RBF-KAN within the PINN framework [Raissi et al., 2019]. To approximate the solution $u$ of a PDE $\mathcal{N}[u]=f$ in domain $\Omega$ subject to boundary conditions $\mathcal{B}[u]=g$ on $\partial\Omega$, we minimize the composite residual loss where derivatives are computed via automatic differentiation. We demonstrate this approach on two benchmarks: 2D heat conduction with high-frequency forcing and the 2D Helmholtz equation with smooth solution.
5.4.1 Heat Conduction in 2D
The heat conduction problem from Wang et al. [2021] has been shown to be more effectively learned by in KAN than by MLP [Wang et al., 2024b]. To evaluate whether the proposed RBF-KAN can also handle multiscale physics-informed problems, we compare RBF-KAN and Free-RBF-KAN against MLP and the original KAN. The governing equation is given by
$u_{t}$ $=\frac{1}{(50\pi)^{2}}u_{xx},$ $\quad(x,t)\in(0,1)\times(0,1),$ (11)
[…]
$u(x,t)$ $=0,$ $\quad(x,t)\in\{0,1\}\times[0,1],$
and its analytical solution is given by $u(x,t)=e^{-t}\sin(50\pi x)$. We randomly sample 4000 interior collocation points and 200 boundary points for each boundary segment. The MLP baseline has 4 hidden layers with width 40, while all KAN variants use an architecture with layer sizes [2,5,5,1] and 30 grid points per activation. The KAN model uses cubic B-spline activations, and the grid range of KAN variations is set as $(x_{l},x_{r})=(0,1)$, same as Wang et al. [2024b]. All models are trained using the Adam optimizer with a learning rate $10^{-3}$ and an exponential learning rate scheduler with $\gamma=0.999$ applied each epoch. Training is performed for 15,000 epochs.
The results presented in Fig. 5 show that both RBF-KAN and Free-RBF-KAN can successfully learn this high frequency problem. While RBF-KAN attains slightly lower accuracy than KAN, it trains in roughly half the time required by KAN. Incorporating adaptive grids and smoothness further improves the performance: Free-RBF-KAN achieves higher accuracy than both KAN and RBF-KAN. Although the additional flexibilities in Free-RBF-KAN increase the training time slighly, Free-RBF-KAN remains substantially faster than KAN, as shown in Table 2.
5.4.2 Helmholtz Equation in 2D
To further demonstrate the benefits of adaptive meshing, we examine physics-informed learning for a 2D Helmholtz equation with Dirichlet boundary conditions, following the setup in Wang et al. [2020]. The problem is defined in the domain $\Omega=(-3,3)^{2}$ as
$\begin{split}u_{xx}+u_{yy}+u&=q\quad\text{ in }\Omega,\\ u&=0\quad\text{ on }\partial\Omega,\end{split}$ (12)
where the forcing term is given to have the analytical solution $u(x,y)=\sin(\pi x)\sin(\pi y)$. For each training epoch, we randomly sampled 4,000 collocation points in the interior and 100 boundary points on each edge of the domain. The MLP baseline has 5 hidden layers with 128 nodes each, whereas all the KAN variants have 2 hidden layers of 5 nodes per layer and 10 grids for each activation function. Training is performed using the Adam optimizer with a learning rate of $10^{-3}$.
Furthermore, the timing results show that KAN is significantly slower than the other approaches, indicating that the B-spline formulation is computationally expensive when combined with AD, as shown by the timings in Table 2. In contrast, RBF-KAN achieves a substantial speed-up.
(a) Exact
(b) KAN
(c) MLP
(d) RBF-KAN
(e) Free-RBF-KAN
Figure 4: The prediction of DeepONet variations on reaction-diffusion PDE. All models has same branch structure with MLP, but with different trunk architecture. A sampled force function ($f$) is used to generate the solution ($u(x,t)$).
5.5 DeepONet with KAN Variants
Our final set of experiments investigates the performance of different architectures within the DeepONet framework. DeepONet [Lu et al., 2019] consists of a branch network and a trunk network; in our study, we evaluate how various choices for the trunk network, including MLP, KAN, RBF-KAN, and Free-RBF-KAN, affect operator-learning accuracy, while the branch network is kept fixed as an MLP. The goal is to learn the solution operator $\mathcal{G}:f(x)\mapsto u(x,t)$ for the 1D nonlinear reaction diffusion equation:
$u_{t}=\frac{1}{100}u_{xx}+\frac{1}{100}u^{2}+f,\quad(x,t)\in(0,1)\times(0,1),$ (13)
with zero boundary and initial condition. All DeepONet variations share the same branch network structure with 100 inputs, 100 outputs, 4 hidden layers of width 40. The trunk net is replaced by MLP, KAN, RBF-KAN, or Free-RBF-KAN, following the configurations summarized in Table 2.
All models were trained with the Adam optimizer using a learning rate of $10^{-3}$, and an exponential learning rate scheduler ($\gamma=0.95$) for 10,000 steps. Training data consist of 50 forcing functions sampled from a Gaussian Random Field (GRF) with the length scale $0.2$, where in each sample 10 locations of sensor observation $u(x,t)$ are randomly selected for training. Performance was evaluated by the relative mean squared error (RMSE) on the full $100\times 100$ spatiotemporal grid, and averaged over 30 random test seeds.
The MLP trunk employs fully connected layers of width 40, yielding the largest model size. In contrast, the KAN-based trunks use a much smaller hidden dimension of 4 and 20 grid points for their basis functions. KAN uses cubic B-splines, whereas the RBF-KAN variants employ Gaussian kernels. The results presented in Table 2 show that Free-RBF-KAN achieves the highest accuracy among all tested architectures while also requiring fewer parameters than the standard KAN.
6 Conclusion
In this paper, we introduced Free-RBF-KAN, an efficient evolution of Kolmogorov–Arnold networks that replaces fixed B-spline bases with adaptive Gaussian kernels. By jointly optimizing grid centroids and shape parameters, our method circumvents the recursive computational overhead of B-splines, achieving a significant reduction in training time on regression and PDE benchmarks while simultaneously improving approximation accuracy. Theoretically, we extended the universal approximation theorem to the RBF-KAN family; empirically, we verified its lack of spectral bias and robust generalization in physics-informed regimes. Although dense MLPs remain the standard for unstructured high-dimensional data, our findings position Free-RBF-KAN as a compelling primitive for scientific computing where interpretability, sparsity, and derivative smoothness are paramount. Future directions include exploring dualities with Gaussian-activated MLPs to leverage mature optimization kernels for further scalability.
Impact Statement
This research contributes to the fundamental advancement of Machine Learning. While such work inherently possesses broad societal implications, we do not identify any specific consequences that require distinct emphasis here.
Appendix
1 Full proof for Eq. 9
@@ 6 Conclusion
2 Solution visualization for experiments in Section 5.4
@@
(k) Free-RBF-KAN
(l) Absolute Error
Figure 5: Predicted solutions from MLP, Free-RBF-KAN, RBF-KAN, and KAN on the heat conduction problem in Section 5.4.1.